A gas storage cylinder in an ordinary chemical labA gas storage cylinder in an ordinary chemical laboratory measures 4.4 cm wide and 17.cm high. This is the label on it.

Contents: N2 gas

Pressure: 29.8 atm

If the cylinder is opened and the gas allowed to escape into a large empty plastic bag, what will be the final volume of nitrogen gas, including what's collected in the plastic bag and what's left over in the cylinder? Write your answer in liters. Round your answer to 2 significant digits.
Answer #1

At constant temperature, assuming equal number of moles,

P1V1 = P2V2

Therefore,

V2 = V1 x (P1 / P2)

Here,

P2 = atmospheric pressure = 1 atm

V2 = volume bag + volume gas left over in cylinder.

Also we have,

V = \pi r2h

Here, Diameter of cylinder = 4.4 cm

Therefore, radius =4.4/2 = 2.2 cm

h =17 cm

Therefore,

V2 = 3.1415 x (2.2)2 x 17 x (1L/1000 cm3) x (29.8/1 atm) = 7.7 L